Accelerator Update for December 10, 2007 - December 12, 2007
Monday December 10
The day shift began with Operations monitoring the Tevatron (TeV) store 5791 (103.06E30) and the Recycler stash (199.1E10), with the Antiproton Source (Pbar) stacking (11.4E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, with NuMI in access, and with SciBooNE / MiniBooNE experiments taking beam.
At 8:36 AM, the High Voltage Coordinator reported a fault on feeder 36, the Proton Fixed Target area. He is investigating.
A MI expert began a coalescing check at 9:56 AM.
NuMI came out of access at 10:39 AM.
The SY sector F1 access ended at 10:42 AM. They didn’t find anything wrong.
Operations established beam to NuMI at 10:58 AM.
The MI coalescing check ended at 11:22 AM.
A SY expert reported at 11:46 AM that they were done with the beam.
Operations prepared Meson enclosures M02 – M05 for access at 1:57 PM.
Pbar experts began a study period at 2:16 PM.
The Neutron Therapy Facility (NTF) reported at 2:21 PM that they had completed treatment on their last patient.
Pbar studies ended at 3:37 PM.
Operations reported at 4:02 PM that the Recycler had stashed antiprotons three times during the day shift.
At 4:03 PM, Operations reported that a Pbar power supply (I:H703) tripped off on a ground fault indication and would not reset. Operations contacted an expert.
A power supply expert reported at 5:01 PM that he had completed repairs on I:H703. A loose hose clamp had allowed LCW to seep into the ground fault detector box.
Pbar resumed stacking at 5:02 PM.
Operations reported at 5:26 PM that the High Voltage Coordinator said the feeder 36 fault was on cable in between PS2 and PS4. He believes that it might be a bad splice. Power had been restored to all buildings.
Operations terminated store 5791 at 6:05 PM.
Operations put the TeV into shot setup at 6:21 PM.
At 6:33 PM, the Linac dump beam vacuum valves closed. Vacuum didn’t recover until 6:53 PM. Linac resumed sending beam.
Operations established store 5792 at 9:11 PM with an initial luminosity of 173.16E30.
At 11:03 PM, a MI power supply (E:QD111) tripped off on a ground fault indication. Operators investigated and then called in a power supply expert.
Operations reported at 11:57 PM that the Recycler had stashed two times during the evening shift.
Tuesday December 11
The midnight shift began with Operations monitoring store 5792 (103.49E30) and the stash (103.6E10), with Pbar stacking (40.6E10), and with NuMI and Sci /MiniBooNE taking beam.
The power supply expert reported that he had replaced a bad SCR on QD111. Pbar resumed stacking and NuMI resumed taking beam at 12:45 AM.
At 1:04 AM, a Linac Klystron RF station (KRF5) tripped off for the first of three times in an hour. It reset.
A Recycler expert brought the Pelletron back online at 7:49 AM.
Operations established beam to SY at 9:29 AM.
At 9:52 AM, Operations reported that the NuMI dehumidifier had tripped off. Operators contacted experts.
At 9:57 AM, Operations reported that LRF4 had tripped off for the first of two times in 30 minutes. Experts investigated, brought it back on, but it tripped off two more times.
After replacing resistors, capacitors, and tightening clamps, Linac experts brought up LRF4 and resumed sending beam at 12:41 PM.
TeV experts began an end-of-store study around 1:43 PM.
Operations reported at 1:54 PM that they had received permission to send beam to the new Meson MTest experiment, T976.
Operations terminated store 5792 at 2:39 PM.
NuMI resumed taking beam at 2:40 PM.
Operations prepared D0 for access at 2:54 PM.
A Sci/MiniBooNE expert reported at 3:30 PM that their autotune program was up and running.
D0 came out of access at 4:10 PM.
Cryo system operators reported at 4:11 PM that they were investigating a problem with the TeV C4 cryo system. They tightened a belt. Operations held off the TeV turn on while the system recovered.
A SY expert reported at 4:35 PM that they had completed their Beryllium target test for T976.
At 4:46 PM, LRF4 tripped off. It reset.
At 4:59 PM, operators had to reset a KRF4 modulator VME timing card.
At 5:02 PM, E:QD111 tripped off on a ground fault indication again. It would not reset. Operations called in a power supply expert.
Operations established beam to the SY absorber at 5:08 PM.
Operations reported at 5:41 PM that beam had been established to the MTest T976 experiment.
Operations put the TeV through a dry squeeze at 6:37 PM.
At 6:53 PM, all the power supplies in the Pbar AP1 line tripped off. Operators investigated.
At 7:04 PM, the TeV sector D1 cold compressor tripped off. Experts came in to make repairs.
A power supply expert reported at 7:39 PM that he had fixed E:QD111. Operations reestablished beam to NuMI.
The Duty Electrician reported at 7:46 PM that he had reset a tripped Pbar yard breaker. He and an operator brought the AP1 line power supplies back online.
Operations put the TeV into shot setup at 10:14 PM. Shot setup halted the MTest beam. The experiment said that they were done for the night.
Wednesday December 12
The midnight shift began with the TeV in shot setup, with Operations monitoring a stash, with Pbar stacking, and with NuMI and Sci/MiniBooNE taking beam.
Operations established store 5795 at 12:32 AM with an initial luminosity of 187.07E30.
At 2:02 AM, a Booster RF station (BRF5) tripped off for the first of several times. Operations bypassed the station and brought BRF19 on in its place.
Power supply experts reported that H717 was back online. Operations performed a critical device test at 3:41 AM and it passed.
Pbar resumed stacking at 342 AM.
At 6:45 AM, Operations turned off the Linac beam in preparation of the scheduled access to repair the Momentum dump vacuum.
A Booster RF expert brought BRF5 back on line at 7:54 AM.
Operations reported at 8 AM that the Recycler had stashed twice during the midnight shift.
The Plans for Wednesday
The plan for today is to complete the Linac Momentum dump vacuum work, which might take up to nine hours. During the work Operations will maintain and monitor the stack, stash, and store.
Comments and Suggestions
|last modified 12/12/2007 email Fermilab|
|Security, Privacy, Legal|